wiki:UC20171009

APO 3.5-m Users Committee Phonecon, 10/09/17

Attending: John Wisniewski (OU), Scott Anderson (UW), Rachel Kuzio de Naray (GSU), John Bally (CU), Mary Beth Kaiser (JHU), Leslie Hebb (NAPG), Mark Klaene (APO), Sarah Tuttle (UW), Kal Kadlec (UW), John Tobin (OU), Anne Verbiscer (UVA; minute taker), Joanne Hughes Clark (Seattle U), Ben Williams (UW)


User feedback and comments from institutional representatives:

  • Wyoming – no one on the call
  • Oklahoma – nothing to report
  • Georgia State – nothing to report
  • Virginia – nothing to report, just lots of snow in Boulder!
  • U. Washington – nothing to report
  • Colorado – nothing to report
  • NAPG - Leslie wanted to know if it’s OK to take dome flats using ARCTIC with mirror covers closed between 18 and 12 degree twilight. The Observing Specialist gave her trouble about it and she didn’t know why. She wants to avoid issues with OS’s in the future. Twilight flats are better for good photometry. Mirror covered dome flats work fine. All depends on level of precision you want later. OS said she shouldn’t be doing dome flats at that time because there would be contamination. Any difference between dome flats and twilight flats will be small. Sounds like a science choice. Only problem staff should have is if it was endangering equipment. Only issues Mark can think of is if you’re closing the mirror covers the telescope goes out of balance. Kills the motors. Also could make staff shift long on winter nights.
  • Seattle – no one on the call
  • Johns Hopkins – nothing to report
  • NMSU - no one on the call

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 9/07/17 through 10/04/17

0) Overview

Rain returned towards the end of the month, hampering observations. ARCSAT resumed observing. DIS and Echelle scatter light being monitored closely. DIS red camera vacuum serviced during a lull.

1) Telescope

Telescope was operational with minimal issues.

2) Instruments

DIS: We started to see scattered light increasing in both red and blue cameras. With a short lull in the schedule it was decided to warm and pump on the red camera. This is now complete and time will confirm how the scattered light behaves in coming couple of weeks.

TripleSpec is operational.

Agile has been operational this period.

Echelle is operational. Inter-order light is higher than we like and may be evidence of CCD or optics contamination. The trend has flat-lined and is not getting worse. Please report if current state is causing issues for your science.

NIC-FPS is operational.

ARCTIC is operational with no change.

3) ARCSAT - The camera control software issues appear to have been mostly cleared, or at least reset back to where it has been with just occasional issues. Survey Camera experienced a loss of vacuum and since it is not user serviceable has been returned to the manufacturer for repair. The time discrepancy issue has been corrected.


Additional telescope and instrument discussion:

There was some additional discussion regarding the increase in scattered light in DIS. The engineering staff did adjust the red camera only because of limited time during this downtime. Blue side is also getting a lot of scattered light. Need to address which side users would prefer. Users prefer to keep scattered light in blue side down. Poll users for their preference. Action item for the committee.

The inter-order light in ARCES is not what we would like it to be but it has stabilized. Users need to reduce their data quickly to see if this is an issue.

ARCSAT has been working but we're having issues with the camera control software. When will SurveyCam be back? Mark says should be 2 weeks, but it had to be sent to Ireland so adding 1 week on each end of that for shipping time (so at least 3 more weeks).


Q4 3.5-m scheduling:

Russet is not on the call. We have filled in a lot of the OPEN time in Q4 schedule. Some remains.


Q4 ARCSAT scheduling:

ARCSAT schedule came out. People are filling in partial weeks. One of the cameras is down.


NA2 Baffle Project

Nancy sent around info. Still waiting for input on what to do. For the most part we know the science drivers. The real question is going to be once we start looking at impact on current instruments on whether it will cause problems. Open to input from users on the science end. Scattered light study is underway. Information gathering is ongoing. John Wilson is not on the call so we cannot get an update from him. How much information does UVa need before they can proceed? Late last week they got information to download. They may need more, but we gave them everything they asked for. What is attenuation at corners of the chip with the current baffling system? We have thrown out numbers at various times but it’s been a while. On the order of 25% is a rough guess. This should flatfield out to a large degree. Unless you’re doing high precision photometry at the corners it should not be a problem. Do you need the high precision over the entire field? The percentage should get added to the report. Joanne joined the call and does do high precision photometry and does need to know the characteristics near the edge of the field. What does she see now? Depends on filters she’s using. Maybe losing 20% with the full sized filters and it depends on the shape of the other filters. She’s willing to help characterize this further. She is losing 20% of photons, not area. She will send this info to Nancy.


TDAWG Update

Had telecon last week. Have nice list of science cases where APO can provide followup on time domain studies. Going well. Will go into report eventually. Next telecon will discuss more of the technical details of the alerts and what they look like, how we access them. Question about whether notes are sent to APO committee. They are not now. Were planning to send final report to committee when it is finished. Shooting for Jan-Feb 2018. Will send to at least the Users Committee for sure.


Report for ARC Board of Governors Meeting

Hoping to get something from NAPG and Johns Hopkins. Update: JHU has sent their package of slides to Nancy.


ACTION ITEMS:

open from previous meetings:

  • Nancy will work with Ben and Russet on a web page that lists the telescope scheduling policies to avoid any future misunderstandings. Status: CLOSED. Nancy, Ben, and Russet added some clarifying text and created the Scheduling procedures and policies page (http://www.apo.nmsu.edu/35m_operations/35m_manual/General_Information/scheduling.html), which is accessible from the main 3.5m scheduling page.
  • Russet will iterate with the UVA team to clarify some of the language in the TripleSpec User's Guide. Status: OPEN. This is an ongoing collaborative effort between Russet and John Wilson, but John has been very busy lately.
  • UC reps should request science highlights from their department members - one Powerpoint slide per program - for sharing at the ARC Board of Governors meeting in early November. The slides can be sent directly to Nancy or collected by the UC reps and sent to Nancy as one packet. Status: CLOSED. Nancy heard from all UC reps.
  • Nancy will distribute a document to the UC reps concerning thoughts on the NA2 baffle project. Status: CLOSED. Nancy sent out the document to the UC on 10/04/17.

new from this meeting:

  • Ben and Nancy will discuss the ARCTIC flats issue on an upcoming 3.5m engineering telecon to figure out what is acceptable for the Observing Specialists, and come up with a policy regarding how late the observers can keep them to take flats.
  • UC reps should poll their DIS users on how they would place priorities on reducing scattered light (red or blue).
  • For the NA2 baffle project, Joanne will send a flat field image around to work with. Based on the information from Joanne Nancy will work with John Wilson to add the amplitude of the current loss of signal in the corners of the ARCTIC FOV to the NA2 Baffle Project Description document.

Next meeting:

The next Users Committee will be on Monday 13 November at 11:00 am MST.

Last modified 10 months ago Last modified on Jul 30, 2021, 7:48:55 AM
Note: See TracWiki for help on using the wiki.